### v1

# Testing stimulated emission photon direction

Abstract
While naively laser only allow to excite external target, for example STED microscopy demonstrates it can also cause its deexitation, however, under uncommon condition of being prepared as excited. These two causations are governed by stimulated emission-absorption pair of equations, and there are presented many arguments that they are switched in perspective of CPT symmetry, however, it would mean photon direction opposite to usually assumed. This article discusses these arguments and proposes simple STED-like tests to experimentally determine this direction once for all, to be performed in a near future. Depending on its result: it should lead to new applications like 2WQC more symmetric quantum computers, or macroscopic violation of CPT symmetry, so far tested only in microscopic settings. Also forward/backward/bidirectional photon sources are discussed.
Keywords: lasers, photonics, STED microscopy, CPT symmetry, ring laser, synchrotron radiation, 2WQC
I Introduction
CPT symmetry (of charge conjugation (C) + parity transformation (P) + time reversal (T)) is shown [1] to be necessary for local Lorentz invariant quantum field theories, hence is generally believed to be satisfied by our nature. It has dozens of successful experimental confirmations [2], however, they are microscopic - come from observation of single particles. Seems conscious macroscopic tests were not performed so far, interesting especially due to T time symmetry which might lead to new applications - some are proposed and discussed here.
As in Fig. 1, applying CPT symmetry to scenario ”laser causes target excitation”, it becomes ”CPT(laser) causes CPT(target) deexcitation”, replacing absorption equation with stimulated emission, with photons traveling from CPT(target) to CPT(laser). In contrast, standard textbook view on stimulated emission is energy being kind of knocked by incoming photon - with the opposite direction for the emitted photon.
As there are at least two possibilities for photon direction, the final judge should be experiment. Related tests were earlier proposed using free electron laser [3], or ring laser [4]. However, they turned out too difficult to organize. This article simplifies test setting to STED-like with just two e.g. diode lasers, determining photon direction through delay. Hence it should be organized in a near future - to update this article and depending on the outcome: leading to further mentioned novel applications, or suggesting macroscopic violation of CPT symmetry.
There are also gathered novel arguments, and proposed view on photon sources: using forward/backward/bidirectional beam as in Fig. 2: applying to external target absorption/stimulated emission for unidirectional, or both equations for bidirectional.
II STED microscopy based tests
Stimulated emission depletion (STED) microscopy [5] improves resolution by reducing size of excited dot with donought-like depletion beam from second laser as in Fig. 2. It uses fluorescent dye being approximately 3 level medium: excitation with one laser from 0 to 2, then fast relaxation to 1, then deexcitation from 1 to 0 with slightly lower wavelength - usually through just spontaneous emission observed by microscope lens.
In STED we additionally speedup this deexcitation through stimulated emission with additional laser of frequency for transition - reducing the number of photons from this region observed by microscope lens focused on spontaneous emission, this way allowing to increase microscope resolution.
While spontaneous emission photons should have a random direction, the question of concern is direction of such stimulated emission photons: toward or outward the laser? Textbooks assume the latter, discussion here based on CPT symmetry suggests the former. There might be also a different angle, what also could be tested experimentally.
II-A Delay STED-like test
Assuming there are only these two options: or , the simplest way to distinguish them seems delay. There is literature discussing delay in STED microscopy, like [6, 7, 8], but they seem insufficient to uniquely answer the direction question.
Therefore, there is proposed STED-like setting as in the bottom of Fig. 1: two (e.g. diode) lasers: continuous for excitation of the dye, and impulse for its stimulated emission. There is also photodetector observing spontaneous emission from this dye, which should observe reduced intensity due to depletion laser impulse - the question is its delay.
Hence photodetector should be connected to oscilloscope, triggered by depletion laser impulse. Its observed delay contains positive contribution for light path from dye to detection, and contribution from laser-dye path, of sign for one of the two assumed possibilities - which could be determined from this delay, e.g. from oscilloscope snapshots planned to be included in the next version of this article.
II-B Alternative STED-like tests
Another approach is placing photodetector in depletion laser-dye line behind the dye (yellow if Fig. 1), allowing to determine or from intensities. Measuring also angle dependence, we could also test (rather exclude) change of angle hypothesis.
As discussed further, additional tests could be using unidirectional e.g. ring laser: its forward beam should not allow for depletion here, its backward beam should allow - what could be also tested in the future.
III CPT symmetry in photonics arguments
Let us gather here some arguments for the expected possibility mainly based on CPT symmetry. Test pointing instead, would need to overcome these arguments, e.g. through macroscopic CPT violation.
III-A Bidirectional and forward/backward beams
Stimulated emission-absorption equations are at heart of lasers, governing deexcitation-excitation e.g. in laser active medium. For atoms in the ground state, excited:
where are (CPT symmetric) Einstein’s coefficients [11], is radiation density of the incident field at frequency , corresponding to transition between the two considered states: .
While both are assumed to act on active medium inside laser, naively only absorption equation acts on external target - for example STED laser demonstrates it is not true, acting e.g. by diode laser with stimulated emission equation on external target. However, this equation is negligible for - requires initial excitation, while unprepared target is close to the ground state.
As in Fig. 2, the situation changes for example for ring laser with optical isolator - using Faraday effect [12] to allow photon circulation only in one direction. Such photons could leave the laser and be absorbed by a (forward) external target.
Looking at such unidirectional ring laser from CPT perspective, photon trajectory would be reversed, causing excitation of different (backward) target - increasing toward minus time, hence decreasing in standard time perspective - as in stimulated emission.
Therefore, this CPT symmetry switches the two equations - in contrast e.g. to diode lasers in STED microscopy, seems only one equation at the time acts on the forward/backward targets - allowing to split them into intuitively forward/backward beam, combined into bidirectional beam in standard e.g. diode laser.
Another example of laser sources switching targets in CPT perspective are based on synchrotron radiation like free electron laser/synchrotron/wiggler/undulator - while there is usually focus only on their forward target, in CPT perspective electron trajectories would be reversed - causing excitation of backward target in the opposite direction, hence should reduce its in standard time perspective, as in stimulated emission.
While application of CPT symmetry seems abstract, general relativity in theory allows a more tangible possibility useful for intuitions - in theory allowing Klein-bottle-like nonorientable wormholes [10] applying P or T symmetry to e.g. a rocket travelling through which - for T symmetry switching absorption/stimulated emission inside, forward/backward beams.
III-B Experimental consequences like Rabi cycle
To test the above with STED-like setting, as mentioned we can replace depletion laser with forward/backward beam - only backward should lead to depletion. If stimulated emission-absorption use opposite photon trajectories, we could also try to split them exploiting delay with fast shutter or switch.
Another type of consequence is population in external target of laser. For bidirectional acting with both equations, we could at most reach population, and it is indeed well known as the final population of damped Rabi cycle [13]: using periodically stimulated emission and absorption equations. In Fig. 3 there is suggested hydrodynamic analog of Rabi cycle (another can be found in [14]), which should exist through EM-hydrodynamics analogy as written, like the shown resonator coupled with external object exchanging momentum in both directions, what might be also appropriate intuition for bidirectional lasers, using cyclically positive and negative radiation pressure [16]. Rabi cycle in CPT perspective should be practically the same - bringing another argument for photons travelling there only between laser and target.
However, using only forward beam, we should be able to reach nearly complete excitation, and such effect has turned out to be known in literature: despite many attempts and theoretical prediction, they were not able to obtain Rabi oscillation with free electron laser [9] - instead getting Autler-Townes effect (dynamics Stark) leading to , instead of for damped Rabi oscillations.
III-C Some possible new applications
If confirmed, the above considerations could lead to many new applications, for example of currently neglected backward beam, like:
Forward beam could allow for state preparation of quantum computers to , backward beam for its CPT analog (postparation) for 2WQC: two-way quantum computers [4].
Backward beam for attacks on quantum cryptographic protocols - use external laser to additionally stimulate emission inside communication laser to steal/intercept and measure such additional photons.
Backward beam for 3D printing/reversed lithography: excite material and precisely stimulate its deexcitation - controlling position and time, also frequency for multiple possible transitions.
Synchrotron sources usually have much higher photon energy, its backward beam could increase probability of some e.g. particle events, which CPT transform use such photons coming from e.g. synchrotron.
IV Conclusions and further work
There were proposed various tests for photon direction of stimulated emission, one of them should be soon performed to update this article. There were also presented many arguments based on CPT symmetry that direction should be from target to laser, which is the opposite than textbook explanation.
The further work depends on the result of the proposed test:
as required by CPT symmetry should lead to many new applications to explore, like the one listed above,
would indicate macroscopic violation of CPT symmetry, requiring further investigation of its details.