### v3

# Testing stimulated emission photon direction

Abstract
While naively laser only causes excitation of external target, for example STED microscopy demonstrates it can also cause its deexitation, however, under uncommon condition of being prepared as excited. These two causalities are governed by stimulated emission-absorption pair of equations, and swap places in perspective of CPT symmetry, however, it means photon direction of stimulated emission should be opposite to usually assumed. This article discusses various arguments and proposes simple STED-like tests to experimentally determine stimulated emission photon direction(s) once for all, possibly both, to be performed in a near future, already having indirect experimental evidence, like lack of Rabi cycle for free electron laser, or recent ”negative time evidence” experiment in the news. Depending on its result: it should lead to new applications like 2WQC more symmetric quantum computers, or macroscopic violation of CPT symmetry, so far probably tested only in microscopic settings. Also proposed forward/backward/bidirectional photon sources are discussed.
Keywords: lasers, photonics, STED microscopy, CPT symmetry, ring laser, synchrotron radiation, 2WQC
I Introduction
CPT symmetry (of charge conjugation (C) + parity transformation (P) + time reversal (T)) is shown [1] to be necessary for local Lorentz invariant quantum field theories, hence is generally believed to be satisfied by our nature. It has dozens of successful experimental confirmations [2], however, it seems all tests so far were microscopic - come from observations of single particles. Conscious macroscopic test was not found in literature, important especially due to T time symmetry which might lead to new applications - some are proposed and discussed here. For example through system response before impulse, like recently experimentally observed [3] in the news, for example in Scientific American 111 www.scientificamerican.com/article/evidence-of-negative-time-found-….
As in Fig. 1, 2, applying CPT symmetry to scenario ”laser causes target excitation”, it becomes ”CPT(laser) causes CPT(target) deexcitation”, replacing absorption equation with stimulated emission, with photons traveling from CPT(target) to CPT(laser). In contrast, standard textbook view on stimulated emission is energy being kind of knocked by incoming photon - with the opposite direction for the emitted photon. As in Fig. 2, the above consequence of CPT symmetry also allows to conclude the textbook view, hence both can be true.
As there are at least two possibilities for photon direction, the final judge should be experiment. Related tests were earlier proposed using free electron laser [4] (earlier 2009 post), or ring laser [5]. However, they turned out too difficult to organize. This article simplifies test setting to STED-like with just two e.g. diode lasers, fluorescent dye and fast photodetector - determining photon direction through delay. Hence it should be organized soon - to update this article and depending on the outcome: leading to further discussed novel applications, or indicating macroscopic violation of CPT symmetry.
There are also gathered novel arguments, and proposed view on photon sources: using forward/backward/bidirectional beam as in Fig. 3: applying to external target absorption/stimulated emission for unidirectional, or both equations for bidirectional.
Updated article version was motivated by ”negative time evidence” in the news [3], which observed clear signal both before and after the impulse, already indicating absorption-stimulated emission causalities of opposite time delays, already bringing some CPT symmetry suggestions. Also, QInter of our 2WQC group (https://www.qaif.org/2wqc): Vysakh Saji has recently performed the proposed experiment, obtaining unofficial result as $\Delta t\approx 0.0082$ ns for $l\approx d$ setting, this way confirming validity of CPT symmetry in macroscopic scale, to be updated with official results. Second update observes that textbook view can be also concluded from CPT symmetry hence both photon directions can be true, also discusses transparency.
Figure 1: Top: recent widely commented experiment by Steinberg group [3] has observed medium response before and after the applied impulse, proposing ”photon can spend a negative amount of time in an atom cloud” type of explanation, which does not seem physical - in contrast to expected here and looking sufficient: stimulated emission. As impulse used in STED microscopy causes both excitation and deexitation of the target, these two causalities/scenarios swap places in CPT symmetry, hence should have opposite photon direction and sign of delay time as observed. Bottom: proposed simplified setting/test, recreating STED situation, but focused on temporal resolution. Continuous laser excites fluorescent dye being approximately 3-level medium, and impulse laser causes its deexcitation through stimulated emission - the question to test is their photon direction, which can be found from $\Delta t=(d\pm l)/c$ delay between laser impulse and observed reduced intensity by fast photodetector focused on dye spontaneous emission. Textbooks assumption would mean ${}^{\prime}+^{\prime}$ sign here, while CPT symmetry requires ${}^{\prime}-^{\prime}$ sign instead, hence this is also macroscopic test of CPT symmetry, complementing dozes of earlier microscopic tests [2].
II STED microscopy based tests
Stimulated emission depletion (STED) microscopy [6] improves resolution by reducing size of excited dot with donought-like depletion beam from second laser as in Fig. 3. It uses fluorescent dye being approximately 3 level medium: excited with one laser from 0 to 2, then fast relaxation to 1, then deexcitation from 1 to 0 with slightly lower wavelength - usually through just spontaneous emission observed by microscope lens.
In STED we additionally speedup this deexcitation through stimulated emission by second: depletion laser of frequency for $1\to 0$ transition - reducing the number of photons from this region monitored by microscope lens focused on spontaneous emission, this way allowing to increase microscope resolution.
While spontaneous emission photons should have a random direction, the question of concern is direction of such stimulated emission photons: toward ()($\leftarrow$)( ← ) or outward ()($\rightarrow$)( → ) the laser? Textbooks assume the latter, discussion here based on CPT symmetry suggests the former. There can be also both, or a different angle, what also could be tested experimentally.
II-A Delay STED-like test
Assuming there are only these two options: ()($\leftarrow$)( ← ) or ()($\rightarrow$)( → ), the simplest way to distinguish them seems delay. There is literature discussing delay in STED microscopy, like [7, 8, 9], however, data they provide seem insufficient to uniquely answer this direction question.
Therefore, there is proposed STED-like setting as in the bottom of Fig. 1: two (e.g. diode) lasers: continuous for $0\to 2$ excitation of the dye, and impulse depletion for its $1\to 0$ stimulated emission. There is also photodetector observing spontaneous emission from this dye, which should observe reduced intensity due to depletion laser impulse - the question is its delay.
Hence photodetector should be connected to oscilloscope, triggered by depletion laser impulse. Its observed delay $\Delta t$ contains positive contribution $d/c$ for light path from dye to detection, and ±$\pm l/c$ contribution from laser-dye path, of sign for one of the two assumed possibilities - which could be determined from this delay, e.g. from oscilloscope snapshots planned to be included in the next version of this article.
Fluorescent dye can be replaced with a different medium for which we can maintain a relatively high population level of excited atoms, what is much more accessible for 3,4-level media. Impulse preferably should have a fraction of nanosecond, (e.g. cascade) photodetector needs to allow nanosecond-scale time resolution. However, their imperfections could be compensated by increasing $l$ distance to impulse laser.
II-B Alternative STED-like tests
Another approach is placing photodetector in depletion laser-dye line behind the dye, allowing to determine ()($\leftarrow$)( ← ) or ()($\rightarrow$)( → ) from intensities. Measuring also angle dependence, we could also test (rather exclude) change of angle hypothesis.
[…]
As discussed further, additional tests could be using unidirectional e.g. ring laser: its forward beam should not allow for depletion here, its backward beam should allow - what could be also tested in the future.
We could also test suggestions of better transparency for backward beam, e.g. by putting (non-excited) obstacle on the way to depletion laser in STED-like setting. It also brings hope for effects on excited targets already for shielded synchrotron sources. Even if such photons could not cross the obstacle, this effect could be used e.g. for information transfer by trying to detect such additional directional photons from excited target behind obstacle, or monitoring its population level.
Rabi cycle could be used instead, by applying impulse laser and monitoring/analyzing situation at impulse boundaries.
III CPT symmetry in photonics arguments
Let us gather here some arguments for the expected ()($\leftarrow$)( ← ) possibility mainly based on CPT symmetry. However, as in Fig. 2, it seems to also imply textbook ()($\rightarrow$)( → ) case, hence both can be true - practically this is test of existence of ()($\leftarrow$)( ← ) effect.
III-A Bidirectional and forward/backward beams
Stimulated emission-absorption equations are at heart of lasers, governing deexcitation-excitation e.g. in laser active medium. For $N_{1}$ atoms in the ground state, $N_{2}$ excited:
[…]
where $B_{12}=B_{21}$ are (CPT symmetric) Einstein’s coefficients [10], $\rho(\nu)$ is radiation density of the incident field at frequency $\nu$, corresponding to transition between the two considered states: $E_{2}-E_{1}=h\nu$.
While both are assumed to act on active medium inside laser, naively only absorption equation acts on external target - for example STED laser demonstrates it is not true, acting e.g. by diode laser with stimulated emission equation on external target. However, this equation is negligible for $N_{2}\approx 0$ requires initial excitation, while unprepared target is close to the ground state.
As in Fig. 3, the situation changes for example for ring laser with optical isolator - using Faraday effect [11] to allow photon circulation only in one direction. Such photons could leave the laser and be absorbed by a (forward) external target.
Looking at such unidirectional ring laser from CPT perspective, photon trajectory would be reversed, causing excitation of different (backward) target - increasing $N_{2}$ toward minus time, hence decreasing $N_{2}$ in standard time perspective - as in stimulated emission. Therefore, this CPT symmetry switches the two equations - in contrast e.g. to diode lasers in STED microscopy, seems only one equation at the time acts on the forward/backward targets - allowing to split beams them into intuitively forward/backward beam acting with separated absorption/stimulated emission equation alone, combined into bidirectional beam in standard e.g. diode laser.
Another example of laser sources switching targets in CPT perspective are based on synchrotron radiation like free electron laser/synchrotron/wiggler/undulator - while there is usually focus only on their forward target, in CPT perspective electron trajectories would be reversed - causing excitation of backward target in the opposite direction, hence should reduce its $N_{2}$ in standard time perspective, as in stimulated emission.
As noticed in Fig. 2, forward/backward beams might have completely different transparency as they act only on $N_{1}$ deexcited / $N_{2}$ excited atoms. As everyday matter usually has $N_{2}/N\approx 0$, backward beam might have much better penetration, allowing for novel applications.
While application of CPT symmetry seems abstract, general relativity in theory allows a more tangible possibility useful for intuitions - in theory allowing Klein-bottle-like nonorientable wormholes [12] applying P or T symmetry to e.g. a rocket travelling through which - for T symmetry switching absorption/stimulated emission inside, forward/backward beams.
Figure 3: Top: ring laser uses optical isolator to enforce unidirectional photon circulation, which would be reversed in CPT perspective - switching the action of absorption and stimulated emission equations, suggesting only one of them should act on external targets - let us refer to them as forward/backward beam. Analogously for synchrotron radiation, wiggler, undulator, free electron laser - reversing electron and photon trajectory in CPT perspective. Center: STED microscopy (image source) showing that, at least for diode lasers, both equations act on external target, hence we can call it bidirectional beam. We usually focus on absorption by external target, because stimulated emission requires its initial excitation, while without such preparation usually $N_{2}/N\approx 0$. In Rabi cycle laser periodically causes excitation and deexitation of target, suggesting requirement of bidirectional beam. Its experimental support is lack of Rabi oscillations for free electron laser [13] - leading to $N_{2}\approx N$ nearly complete excitation in Autler-Townes effect (dynamic Stark), instead of $N_{2}\approx N/2$ for damped Rabi cycle. Bottom: while CPT symmetry of physics is required by local Lorentz-invariant QFT, e.g. switching forward/backward beams, in theory general relativity allows its realization by travelling through nonorientable wormhole [12], bringing tangible thought experiment valuable for intuitions.
III-B Experimental consequences like Rabi cycle
To test the above with STED-like setting, as mentioned we can replace depletion laser with forward/backward beam - only backward should lead to depletion. If stimulated emission-absorption use opposite photon trajectories, we could also try to split them exploiting delay with fast shutter or switch, or maybe exploiting transparency absorbing only forward beam.
Another type of consequence is $N_{1}-N_{2}$ population in external target of laser. For bidirectional acting with both equations, we could at most reach 1/21/$2-1/$ 2 - 1 / 2 population, and it is indeed well known as the final population of damped Rabi cycle [14]: using periodically stimulated emission and absorption equations. In Fig. 4 there is suggested hydrodynamic analog of Rabi cycle (another can be found in [15]), which should exist through EM-hydrodynamics analogy as written, like the shown resonator coupled with external object exchanging momentum in both directions, what might be also appropriate intuition for bidirectional lasers, using cyclically positive and negative radiation pressure [16]. Rabi cycle in CPT perspective should be practically the same - bringing another argument for photons travelling there only between laser and target.
However, using only forward beam, we should be able to reach $N_{2}\approx N$ nearly complete excitation, and such effect has turned out to be known in literature: despite many attempts and theoretical prediction, they were not able to obtain Rabi oscillation with free electron laser [13] - instead getting Autler-Townes effect (dynamics Stark) leading to $N_{2}\approx N$, instead of $N_{2}\approx N/2$ for damped Rabi oscillations.
III-C Some possible new applications
If confirmed, the above considerations could lead to many new applications, especially of currently neglected backward beam, like:
  • Forward beam (e.g. by adding optical isolator) could allow for state preparation of quantum computers to nearly completely excited $N_{2}\approx N$ as in [13], backward beam for its CPT analog (postparation) for 2WQC: two-way quantum computers [5] in theory allowing to solve NP problems [17].
  • If we could send information this way, there could be e.g. built time-loop computers (Section V of [4]) for example using backward beam to send chip output back to its original input with negative time delay, also in theory allowing to solve NP problems. Connecting multiple such negative time transmitters e.g. into a cycle, could allow to also reach macroscopic negative time differences.
  • Backward camera e.g. exploiting transparency or for astronomical observations - focused on target’s backward beam, while standard camera is focused on its forward beam. It would replace CMOS/CCD matrix behind the lens, with matrix of continuously excited e.g. fluorescent dye, plus photodetectors monitoring their spontaneous emission.
  • Backward beam should be transparent for $N_{2}=0$ medium, what might allow for much better penetration than forward (transparent for $N_{1}=0$) - for example for backward camera or communication through obstacles.
  • Backward beam for attacks on quantum cryptographic protocols like BB84 [18] - use external laser to additionally stimulate emission inside communication laser to steal/intercept such additional photons, which could be read e.g. by passing through control of CNOT gate - affecting second qubit to be measured.
  • Backward beam for observations actively increasing deexcitation rate of excited objects, e.g. for thermal imaging,
  • Backward beam for 3D printing/reversed lithography: excite material and precisely stimulate its deexcitation - controlling position and time, also frequency to choose one of multiple possible transitions, e.g. some high energy transition to locally reduce temperature.
  • Synchrotron sources usually have much higher photon energy, its backward beam could e.g. increase probability of selected nuclear decays/transitions: producing photons of chosen energy. Maybe also for other particle events: which CPT transform use such photons coming from e.g. synchrotron.
IV Conclusions and further work
There were proposed various tests for photon direction of stimulated emission, soon to be performed to update this article. There were also presented many arguments based on CPT symmetry that direction should be from target to laser, which is the opposite than textbook explanation, but allowing to imply it as a consequence. Analogous processes could be also possible for other EM frequencies like radio to be explored.
The further work depends on the result of the proposed test:
  • ()($\leftarrow$)( ← ) as required by CPT symmetry should lead to many new applications to explore, like the one listed above,
  • lack of ()($\leftarrow$)( ← ) would indicate macroscopic violation of CPT symmetry, requiring further investigations of its details.
Figure 4: Top: Mathematical analogy between electrodynamics and hydrodynamics (superfluid - no viscosity) from [19]. Center: therefore, Rabi cycle also needs hydrodynamical analog, tank of liquid with standing wave seems analogous to laser resonator. Being semi-transparent/semi-open in one direction, it should cause left-right swimming of a probe particle outside, through exchange of momentum acting in both directions. In CPT perspective it looks the same, also with momentum exchange in both directions. Bottom: while the above bidirectional source is resonator kind of extending to external target e.g. for Rabi oscillations, there are also unidirectional sources. In this analogy, synchrotron radiation is kind of water stream allowing only unidirectional momentum transfer. Optical isolator analog could be some unidirectional valve, analog of ring laser a pump enforcing unidirectional circulation - with reversed circulation in CPT perspective. Radiation pressure is a vector: can be also negative [16].