### v6

# Testing stimulated emission photon directions

Abstract
While naively laser only causes excitation of external target, e.g. Rabi cycle, STED microscopy or backward ASE in optical amplifiers demonstrate it can also cause its deexitation, however, under uncommon condition of being prepared as excited. These two causalities are governed by absorption-stimulated emission pair of equations, and swap places in perspective of CPT symmetry, however, it means photon direction of stimulated emission should be opposite to usually assumed. This article discusses various arguments and proposes simple direct tests to experimentally verify existence of such backward photon trajectories, complementing consequent forward textbook trajectories. There are already indirect experimental arguments, like backward ASE, lack of Rabi cycle for free electron laser, or recent ”negative time evidence” experiment in the news. Depending on the result: it should lead to new applications like 2WQC more symmetric quantum computers, or macroscopic violation of CPT symmetry, so far probably tested only in microscopic settings. Also proposed forward/backward/bidirectional photon sources are discussed.
Keywords: lasers, photonics, STED microscopy, backward ASE, CPT symmetry, ring laser, synchrotron radiation, 2WQC
I Introduction
Figure 1: Top: recent widely commented experiment by Steinberg group [1] has observed medium response (measured phase of 85Rb) before and after the applied impulse, proposing ”photon can spend a negative amount of time in an atom cloud” type of explanation, which does not seem physical.
Bottom: instead, as discussed here, it might be a result of just stimulated emission due to EM coupling of laser resonator and atoms as tiny resonators, required as some of the atoms were excited by magneto-optical trap (MOT) lasers, and if being CPT analog of absorption, should have opposite delay sign.
CPT symmetry (of charge conjugation (C) + parity transformation (P) + time reversal (T)) is shown [2] to be necessary for local Lorentz invariant quantum field theories, hence is generally believed to be satisfied by our nature. It has dozens of successful experimental confirmations [3], however, it seems all tests so far were microscopic - come from observations of single particles. Conscious macroscopic test was not found in literature, important especially due to T time symmetry which might lead to new applications - some are proposed and discussed here. For example through system response before impulse, like recently experimentally observed by Steinberg group in the news [1]111 www.scientificamerican.com/article/evidence-of-negative-time-found-….
Figure 2: Top: CPT symmetry says that if there exists ”laser causes excitation” scenario, also ”laser causes deexciation” should exist. Being CPT symmetry analogs, they should have opposite photon direction and delay sign - we would like to verify it experimentally and apply if successful. For example in STED microscopy diode lasers cause both.
Bottom: proposed STED-like test of this photon direction(s): continuous laser excites fluorescent dye being approximately 3-level medium, and impulse laser causes its deexcitation through stimulated emission - its photon direction can be found from $\Delta t=(d\pm l)/c$ delay between laser impulse and observed reduced intensity by fast photodetector focused on dye spontaneous emission. Textbooks assumption would mean ${}^{\prime}+^{\prime}$ sign here, while CPT symmetry requires ${}^{\prime}-^{\prime}$ sign instead, hence this is also macroscopic test of CPT symmetry, complementing dozes of earlier microscopic tests [3].
Figure 3: In commercially available e.g. erbium-doped optical fiber amplifiers (EDFA, or simpler semiconductor SOA, etc.) [4] there is observed strong backward ASE (amplified spontaneous emission) - increased emission from pumped doped fiber toward laser, exactly as in the sought CPT analog of ”laser causes excitation” effect. Top:Analogously to STED-like test, we can verify this photon direction by measuring delay between the impulse and reduced population of excited atoms of external target, e.g. by monitoring its spontaneous emission. For this purpose, the shown standard forward optical isolator should be removed, or better rotated to backward position. As erbium has very long millisecond-scale lifetimes, there is required fast photodiode. Fig. 7 contains further simplified test. Bottom: simplified universal setting as in [5], more practical to measure backward ASE delay, separation, transparency, and for their potential applications.
As in Fig. 2, applying CPT symmetry to ”laser causes target excitation” scenario, it becomes ”CPT(laser) causes CPT(target) deexcitation”, replacing absorption equation with stimulated emission, having photons traveling from CPT(target) to CPT(laser), also reversing time delay, what might explain observations of negative time delays, also e.g. [6], finally being a consequence of large number of single emissions/absorptions.
We would like to verify existence of such CPT analog of ”laser causes excitation” scenario with backward photon trajectory, for example in STED-like setting as in Fig. 2. Well known backward ASE in common optical amplifiers, as in Fig. 3, seems exactly the sought scenario, also providing practical realization of such test, and maybe also discussed further potential applications thanks to wide commercial availability of optical fiber hardware.
In contrast, as in Fig. 4, standard textbook view on stimulated emission is energy being kind of knocked forward by incoming photon - with the opposite direction for the emitted photon. There are lots of experimental evidence for existence of such process. While it seems in disagreement with CPT symmetry, bottom of Fig. 4 shows it can be viewed as a consequence of the sought: CPT analog of ”laser causes excitation” scenario.
As there are at least these two possibilities for stimulated emission photon directions, the final judge should be experiment. Related tests were earlier proposed using free electron laser [7] (earlier 2009 post), or ring laser [8]. However, they turned out too difficult to organize. This article simplifies test setting to STED-like, and later to optical amplifier - determining photon direction through delay. Hence it should be organized soon - to update this article, and depending on the outcome: leading to further discussed novel applications, or indicating macroscopic violation of CPT symmetry.
Absorption/stimulated emission equations are switched in CPT perspective, and e.g. diode laser in STED microscope acts with both on external target (bidirectional beam), also Rabi cycle requires both. There is a natural question if we can separate them - there will be shown arguments that e.g. optical isolator or synchrotron radiation allows to split them - into forward/backward unidirectional beam as in Fig. 6.
Multiple updates of this article were mainly motivated by finding additional indirect experimental evidence in the literature (e.g. for Fig. 1, 3, 5), transparency with textbook view as consequent, convenient optical amplifier realization, and adding multiple potential applications. Also, QInter of our 2WQC group (https://www.qaif.org/2wqc): Vysakh Saji has recently performed the proposed experiment, obtaining unofficial result as $\Delta t\approx 0.0082$ ns for $l\approx d$ setting, this way confirming validity of CPT symmetry in macroscopic scale, to be updated with official results.
[…]
Center: basic argument for ()($\leftarrow$)( ← ) is CPT symmetry of physics: transforming ”laser causes target excitation” scenario into ”CPT(laser) causes CPT(target) deexcitation”, also switching the stimulated emission-absorption equations, governed by the same (CPT symmetric) Einstein’s coefficients $B_{12}=B_{21}$.
Bottom: schematic explanation that CPT view ()($\leftarrow$)( ← ) also allows to conclude textbook ()($\rightarrow$)( → ) picture, hence both can be true. For photons travelling in ”laser causes excitation/deexcitation” CPT analogous scenarios, we add a medium on the way: looking at transparency in the latter, stimulated emission should fully act on the target, but some photons it produces should be absorbed by medium - returning from CPT perspective it becomes textbook amplification by going through excited medium. All should be coupled by standing waves as in Fig. 7, stimulating/amplifying all these photon emissions.
Figure 5: Other type of argument for time/CPT symmetry of lasers are pre-pulses generated by post-pulses, observed (e.g. [9, 10]) for extremely high energy lasers based on chirped-pulse amplification (CPA) going through nonlinear medium (being first stretched then compressed). Theoretical analysis is based on Fourier transform in time, being symmetrized by nonlinearities. However, decomposing it into individual absorptions/emissions, negative delays should again come from being CPT analogs, hence having opposite delay signs.
II STED microscopy based tests
Stimulated emission depletion (STED) microscopy [11] improves resolution by reducing size of excited dot with donought-like depletion beam from second laser as in Fig. 6. It uses fluorescent dye being approximately 3 level medium: excited with one laser from 0 to 2, then fast relaxation to 1, then deexcitation from 1 to 0 with slightly longer wavelength - usually through just spontaneous emission observed by microscope lens.
In STED we additionally speedup this deexcitation through stimulated emission by second: depletion laser of frequency for $1\to 0$ transition - actively reducing the number of photons from doughnut-shaped region. This way the final excited dot is smaller, increasing microscope resolution for lens focused on stimulated emission of the dye.
While spontaneous emission photons should have a random direction, the question of concern is direction of stimulated emission photons: ()($\rightarrow$)( → ) forward ano/or ()($\leftarrow$)( ← ) backward? Textbooks assume forward, discussion here based on CPT symmetry suggests backward as fundamental, and forward as its consequence (Fig. 4). Therefore, both can be true, the question to verify is existence of backward ()($\leftarrow$)( ← ).
II-A Delay STED-like test
Assuming there are only these two options: ()($\leftarrow$)( ← ) or ()($\rightarrow$)( → ) and we would like to test existence of the former, the simplest way to distinguish them seems delay. There is literature discussing delay in STED microscopy, like [12, 13, 14], however, data they provide seems insufficient to uniquely answer this question.
Therefore, there is proposed STED-like setting as in the bottom of Fig. 2: two (e.g. diode) lasers: continuous for $0\to 2$ excitation of the dye, and impulse depletion for its $1\to 0$ stimulated emission. There is also photodetector observing spontaneous emission from this dye, which should observe reduced intensity due to depletion laser impulse - the question is its delay.
Hence (fast) photodetector should be connected to oscilloscope, triggered by depletion laser impulse. Its observed delay $\Delta t$ contains positive contribution $d/c$ for light path from dye to detector, and ±$\pm l/c$ contribution from impulse laser-dye path, of sign for one of the two assumed possibilities - which could be determined from this delay, e.g. from oscilloscope snapshots planned to be included in the next version of this article.
Fluorescent dye can be replaced with a different medium for which we can maintain a relatively high population level of excited atoms, what is much more accessible for 3,4-level media. Impulse preferably should have a fraction of nanosecond, (e.g. cascade) photodetector needs to allow nanosecond-scale time resolution. However, their imperfections could be compensated by increasing $l$ distance to impulse laser.
As in Fig. 3, standard commercially available EDFA fiber optics equipment seems very convenient for analogous test and further applications if successful (suggested by Michał Markiewicz). The sought target $\to$ laser photon trajectories caused by laser seems exactly the backward ASE effect, well known in literature - usually unwanted and removed by optical isolator. We can test these photon trajectories by removing or better reversing this optical isolator, adding (rather fast) photodetector filtered to observed the stimulated emission, and observing delay of its reduction as response for impulse of deexcitation laser. In comparison to STED setting, $d/c$ propagation time is between EDF and detector, $l/c$ between EDF and impulse laser - easily extended by delay line.
II-B Alternative STED-like tests
Another approach is placing photodetector in depletion laser-dye line behind the dye, allowing to determine ()($\leftarrow$)( ← ) or ()($\rightarrow$)( → ) from intensities. Measuring also angle dependence, we could also test (rather exclude) change of angle hypothesis.
As discussed further, additional tests could use unidirectional e.g. ring laser (or standard with added optical isolator): its forward beam should not allow for depletion here, its backward beam should allow - what could be also tested in the future.
We could also test suggestions of better transparency for backward beam, e.g. by placing (non-excited) obstacle on the way to depletion laser in STED-like setting. It also brings hope for effects on excited targets already for shielded synchrotron sources. Even if such photons could not cross the obstacle, this effect could be used e.g. for information transfer by trying to detect such additional directional photons from excited target behind obstacle, or monitoring its population level.
Rabi cycle could be used instead, by applying impulse laser and monitoring/analyzing situation at impulse boundaries, or including optical isolator on the way.
Figure 6: Top: ring laser uses optical isolator to enforce unidirectional photon circulation, which would be reversed in CPT perspective - switching the action of absorption and stimulated emission equations, suggesting only one of them should act on external targets - let us refer to them as forward/backward beam. Analogously for just adding optical isolator, or synchrotron sources like wiggler, undulator, free electron laser - reversing electron and photon trajectory in CPT perspective.
Center: STED microscopy (image source) showing that, at least for diode lasers, both equations act on external target, hence we can call it bidirectional beam. We usually focus on absorption by external target, because stimulated emission requires its initial excitation, while without such preparation usually $N_{2}/N\approx 0$. In Rabi cycle laser periodically causes excitation and deexitation of target, suggesting requirement of bidirectional beam. Its experimental support is lack of Rabi oscillations for free electron laser [15] - leading to $N_{2}\approx N$ nearly complete excitation in Autler-Townes effect (dynamic Stark), instead of $N_{2}\approx N/2$ for damped Rabi cycle.
Bottom: while CPT symmetry of physics is required by local Lorentz-invariant QFT, e.g. switching forward/backward beams, in theory general relativity allows its realization by travelling through nonorientable wormhole [16], bringing tangible thought experiment valuable for intuitions.
III CPT symmetry in photonics arguments
Let us gather here some arguments for the expected ()($\leftarrow$)( ← ) possibility mainly based on CPT symmetry. However, as in Fig. 4, it seems to also imply textbook ()($\rightarrow$)( → ) case, hence both can be true - hence practically this is a test of existence of ()($\leftarrow$)( ← ) effect.
III-A Bidirectional and forward/backward beams
Stimulated emission-absorption equations are at heart of lasers, governing deexcitation-excitation e.g. in laser active medium. For $N_{1}$ atoms in the ground state, $N_{2}$ excited:
[…]
where $B_{12}=B_{21}$ are (CPT symmetric) Einstein’s coefficients [17], $\rho(\nu)$ is radiation density of the incident field at frequency $\nu$, corresponding to transition between the two considered states: $E_{2}-E_{1}=h\nu$.
While both are assumed to act on active medium inside laser, naively only absorption equation acts on external target - for example STED laser demonstrates it is not true, acting e.g. by diode laser with stimulated emission equation on external target. However, this equation is negligible for $N_{2}\approx 0$ requires initial excitation, while unprepared target is close to the ground state.
As in Fig. 6, the situation changes for example for ring laser with optical isolator - using Faraday effect [18] to allow photon circulation only in one direction. Such photons could leave the laser and be absorbed by a (forward) external target.
Looking at such unidirectional ring laser from CPT perspective, photon trajectory would be reversed, causing excitation of different (backward) target - increasing $N_{2}$ toward minus time, hence decreasing $N_{2}$ in standard time perspective - as in stimulated emission. Therefore, this CPT symmetry switches the two equations - in contrast e.g. to diode lasers in STED microscopy, seems only one equation at the time acts on the forward/backward targets - allowing to split beams into intuitively forward/backward beam acting with separated absorption/stimulated emission equation alone, combined into bidirectional beam in standard e.g. diode laser.
Another example of laser sources switching targets in CPT perspective are based on synchrotron radiation like free electron laser/synchrotron/wiggler/undulator - while there is usually focus only on their forward target, in CPT perspective electron trajectories would be reversed - causing excitation of backward target in the opposite direction, hence should reduce its $N_{2}$ in standard time perspective, as in stimulated emission.
As noticed in Fig. 4, forward/backward beams might have completely different transparency as they act only on $N_{1}$ deexcited / $N_{2}$ excited atoms. In other words, while textbook view on stimulated emission means amplification of forward beam by going through excited medium, its CPT analog is amplification of backward beam by going through deexcited medium (as in this Figure). Therefore, as everyday matter usually has $N_{2}/N\approx 0$, backward beam might have much better penetration, allowing for novel applications. However, coupling standing wave might need to reach the target, what could essentially weaken such effect.
While application of CPT symmetry seems abstract, general relativity in theory allows a more tangible possibility useful for intuitions - in theory allowing Klein-bottle-like nonorientable wormholes [16] applying P or T symmetry to e.g. a rocket travelling through which - for T symmetry switching absorption/stimulated emission inside, forward/backward beams.
III-B Optical cavity and delocalized standing EM wave coupling
In vacuum Rabi cycle [19] e.g. a single atom periodically exchanges energy with optical cavity - cycling energy between atomic orbital, and standing EM wave of cavity, with energy density more localized in the former case. This difference is similar to between laser and radio communication - the former uses localized optical photons travelling in one direction, the latter delocalized e.g. spherical EM waves travelling in all directions.
While both are usually described with Fock space, this is abstract effective description of pertubative approximation. Like ”apple + apple = 2 apples” correct algebra, for deeper dependencies requiring to ask about structure, field configuration in non-perturbative picture, here mainly of electromagnetic fields - again with higher energy density when localized in atom e.g. during vacuum Rabi oscillations.
Being different types of EM field configurations, their interactions might also differ, for example with sophisticated devices like optical isolator - specialized to block localized optical photons travelling in one direction, but not necessarily delocalized standing EM waves for coupling - what should be tested, e.g. with setting as in Fig. 7. Finding material which acts as a mirror for optical photons, but is semi-transparent for delocalized EM waves for coupling, would be perfect to close resonator of potential backward lasers: only amplifying target deexcitation, e.g. for radiotherapy as in Fig. 10.
III-C Rabi cycle as coupling, its asymmetrization
Generally, in Rabi cycle as in Fig. 7, analogously to coupled pendula (e.g. [20, 21, 22, 23]), there is observed coupling of two resonators periodically exchanging energy and momentum in both directions - in practice using mainly lasers as pumped directional resonators, atoms as tiny isotropic resonators, or cavities leading to standing waves/eigenmodes.
As electromagnetism and hydrodynamics are governed by nearly the same equations, it is also valuable to think about Rabi cycle from hydrodynamics perspective as in Fig. 8: laser as standing wave extending outside resonator due to semi-transparent wall, exchanging energy and momentum with external objects in both directions - with positive and negative pressure, both exist also for radiation [24].
Such bidirectional energy exchange for e.g. laser-atoms coupling is mainly realized by optical photons, hence we can view standard e.g. diode laser as bidirectional beam - causing both excitation of target, but also deexcitation - e.g. in Rabi cycle, STED microscopy, or backward ASE of EDFA.
Being able to introduce asymmetry in this coupling could lead to new applications. For example free electron laser (FEL) does not have such resonator coupling, hence instead of damped Rabi [25] leading to $N_{2}\to N/2$, FEL experiments [15] show $N_{2}\to N$ only excitation with unidirectional: forward beam - photon exchange only in one direction, acting only with absorption equation on the target. CPT symmetry suggests that looking at such synchrotron radiation beam from CPT perspective, photon beam created by reversed electron trajectory should act as backward beam - causing only target deexctiation, acting only with stimulated emission equation.
As synchrotron radiation sources are large and costly, for practical reasons we should search ways to separate absorption and stimulated emission equations with different tools like optical isolator, ring laser - to be tested.
Figure 7: Top:coupled pendula (image source) as analog of Rabi cycle (e.g. [20, 21, 22, 23]) - to view it as coupling of e.g. laser as pumped directional resonator, and atom as tiny isotropic resonator, or standing waves of optical cavity as resonator e.g. through Jaynes-Cummings model [26].
Center: vacuum Rabi cycle [19] - observed oscillations between optical cavity resonator, and atom as tiny resonator - of energy between being localized in atom (high energy density), and delocalized standing EM wave (low energy density) - having different EM fields, like localized laser vs delocalized radio communication. Laser can enforce Rabi oscillations, periodically causing target excitation and deexcitation - allowing to view these causalities as results of coupling of two resonators, in contrast e.g. to free electron laser without resonator - hence only observing target excitations there [15]: $(N_{2}\to N)$ instead of damped Rabi $(N_{2}\to N/2)$. Being able to add some asymmetry to such energy transfer caused by coupling, as discussed further can lead to many applications - e.g. optical isolator brings hope for that: designed to block localized optical photons in one direction, what not necessarily has to be true for delocalized standing wave for coupling - having lower energy density.
Bottom: proposed test if optical isolator blocks this EM coupling - monitoring backward ASE: amplified emission from erbium toward laser, for 3 optical isolator cases. It is observed without optical isolator (’none’ case), backward ASE photons are often blocked by $\rightarrow$ forward placed optical isolator. The main question is if $\leftarrow$ backward placed optical isolator leaves backward ASE - suggested both by CPT symmetry (switching photon and optical isolator direction), and coupling through delocalized EM wave.
Figure 8: Top: Mathematical analogy between electrodynamics and hydrodynamics (superfluid - no viscosity) from [27].
Center: therefore, Rabi cycle also needs hydrodynamical analog, tank of liquid with standing wave seems analogous to laser resonator. Being semi-transparent/semi-open in one direction (also in Fig. 7), it should cause left-right swimming of a probe particle outside, through exchange of momentum acting in both directions, using positive/negative pressure. In CPT perspective it looks the same, also with energy/momentum exchange in both directions.
Bottom: while the above bidirectional source is resonator extending to external targets, coupling with them through standing wave e.g. for Rabi oscillations, there are also unidirectional sources. In this analogy, synchrotron radiation is kind of water stream allowing only unidirectional momentum transfer. Optical isolator analog could be some unidirectional valve, analog of ring laser: a pump enforcing unidirectional circulation - with reversed circulation in CPT perspective. Radiation pressure is a vector: can be also negative [24], suggesting to build more symmetric and powerful two-way microfluidic, photonic computers as in Fig. 9.
IV Some potential new applications
If confirmed, the above considerations could lead to many new applications, especially for the currently neglected backward beam - optimistically of standard e.g. optical fiber equipment (resonator coupling) or ring lasers, pessimistically of synchrotron radiation (forward/backward sources), like:
  • Forward beam, e.g. from synchrotron radiation, could allow for state preparation of quantum computers to nearly completely excited $N_{2}\approx N$ as in [15], backward beam for its CPT analog (postparation) for 2WQC [8] as in Fig. 9 in theory allowing for up to exponential speedup and better stability ([28, 29]).
  • If we could send information this way, there could be e.g. built time-loop computers (Section V of [7]) for example using backward beam to send output of chip back to its original input with negative time delay, also in theory allowing to solve NP problems. Connecting such multiple negative delay transmitters e.g. into a cycle, could allow to also reach macroscopic negative time differences.
  • Backward beam might be transparent for $N_{2}=0$ medium, what could allow for much better penetration than forward (transparent for $N_{1}=0$) - e.g. for backward camera as in Fig. 11, various scanners, communication through obstacles. Finding material being mirror for optical photons, but semi-transparent for standing EM waves for coupling, it would be perfect to close resonator of potential backward laser.
  • Backward beam to stimulate specific chemical processes: releasing photons of chosen energy (and polarization), e.g. for chemical/technological processes, energy production, or medical therapies as in Fig. 10 - e.g. locally reduce lifetime of ATP molecule (stimulate decay also releasing this energy locally heating), or inhibit biochemical processes by amplifying deexcitation of crucial molecules, for example with multiple beams intersecting in cancer tissue, especially if better transparency is confirmed.
  • Backward camera maybe exploiting transparency and/or opposite delay e.g. for astronomical observations - focused on target’s backward beam, while standard camera is focused on its forward beam. It would replace CMOS/CCD matrix behind the lens, with matrix/surface of continuously excited approximately 3-level medium like dye, plus some monitoring of its population map - e.g. through spontaneous emission, or EIT (electromagnetically induced transparency).
  • Backward beam for attacks on quantum cryptographic protocols like BB84 [30] - use external laser to additionally stimulate emission inside communication laser (as backward ASE) to steal/intercept such additional photons, which could be read e.g. by passing through control qubit of CNOT gate - affecting the second qubit to be measured.
  • Backward beam for observations actively increasing deexcitation rate of excited objects, e.g. for thermal imaging.
  • Backward beam for 3D printing/reversed lithography: excite material and precisely stimulate deexcitation - controlling position, time, polarization, also frequency to choose one of multiple possible transitions e.g. of high energy to locally reduce temperature (cooling), of energy regained thermally.
  • Synchrotron sources usually have much higher photon energy, its backward beam could e.g. increase probability of selected nuclear decays/transitions: producing photons of chosen energy, polarization. Maybe also for other particle events: which CPT transform requires such photons incoming from e.g. synchrotron.
Figure 9: Original motivation: (example of) two-way quantum computer (2WQC) [8] as adding a process which in CPT perspective becomes the original state preparation process. For photonic QC state preparation is usually performed by laser impulse, suggesting to just add ’reversed’ impulse: delayed by propagation time, and behind backward placed optical isolator - this way from CPT perspective photon trajectories should be reversed. In comparison to standard 1WQC, in theory 2WQC allows for up to exponential speedup e.g. to solve NP problems, also better stability and error correction [28, 29]. Such classical 2WCC photonic computer on continuous beam might be also interesting.
Figure 10: Being able to split absorption and stimulated emission equations, for example with CPT analog of synchrotron radiation: beam resulting from reversed electron trajectories as in Fig. 6, or maybe standard laser with backward placed optical isolator as in Fig. 7, one of many new applications could be radiotherapy trying to locally amplify emission of photons of chosen energy and polarization. For example to try to locally increase degradation of energy carrying molecules like ATP (if having emission spectrum) - to starve cancer tissue, and heat up by releasing this energy. Intersecting multiple such beams, maybe additionally rotating them for fixed intersection point, we could make the effect there orders of magnitude stronger then for intermediate tissues, to make it negligible outside of targeted tissue. Different issues could be also targeted this way, e.g. trying to disable toxic molecules, inhibit chosen biochemical pathways by stimulating deexctiation of crucial molecules e.g. of pathogens, or just for local cooling, etc.
Figure 11: Top: Backward camera observing backward ASE caused by the target, especially interesting if confirmed better transparency and/or opposite delay sign. Can be e.g. based on (multi-mode) fiber with circulators as in [5] and Fig. 3, or more practical with semitransparent mirror and CCD/CMOS - both need some continuously pumped medium (maybe from some display technology). Alternatively one could monitor spontaneous emission of the pumped medium.
Bottom: Backward lighted camera: applying backward ASE effect on the target to amplify its emission in chosen spectrum toward camera (e.g. thermal also cooling), or on laser medium used to send quantum cryptography protocols like BB84, to increase its number of produced photons and steal/read some.
V Conclusions and further work
There were proposed various tests for photon direction of stimulated emission, soon to be performed to update this article. CPT symmetry requires additional to textbook direction, to be verified. Analogous processes could be also possible for other EM frequencies like radio, what is worth exploring.
The further work depends on the result of the proposed test:
  • ()($\leftarrow$)( ← ) as required by CPT symmetry should lead to many new applications to explore, like the one listed above,
  • lack of ()($\leftarrow$)( ← ) would indicate macroscopic violation of CPT symmetry, requiring further investigations of its details.